# Hierarchical Refinement of Universal Multimodal Attacks on Vision-Language Models

### v1

@@ IV-B2 Effects of multimodal attacks
Incorporating text attacks further enhances the attack performance, as evidenced by clear gains of C-PGC and HRA over other methods. Replacing the most salient words yields additional improvements, which is reflected in the superior performance of HRAimp compared with HRAran. Interestingly, although C-PGC also replaces the most important words, our HRAran still performs better in many cases, indicating the effectiveness of directly discovering substitution words rather than learning universal embeddings and then mapping them back to tokens. Moreover, our method does not rely on a predefined word library, further demonstrating its practicality. It is also noted that identifying the most important words would incur additional computational overhead (i.e., HRAimp), while our HRAran avoids.
@@ IV-B3 Effects of learning objectives
Learning objectives also influence transferability. Models trained with similar objectives tend to exhibit stronger adversarial transferability, as observed in attacks across different CLIP variants and other architectures. For instance, UAPs generated on one CLIP model transfer better to other CLIP models than to ALBEF or TCL.
@@ IV-B4 Asymmetry of attacks
These results and analyses demonstrate that adversarial transferability is shaped by multiple factors, including data properties, model architectures, learning objectives, and cross-modal interactions. This highlights the importance of accounting for these factors when designing transferable attacks.
@@ IV-C1 From image-text retrieval to other tasks
We first examine transfer attacks from image–text retrieval to visual grounding and image captioning, with results presented in Tables IV and III, respectively. For both experiments, we generate UAPs using CLIP and ALBEF as the source models and Flickr30K as the source dataset. For visual grounding, we use ALBEF and RefCOCO+ as the target model and dataset. For image captioning, we use BLIP and MSCOCO as the target model and dataset.
Across both tasks, all methods exhibit notably reduced transferability. This is expected due to substantial differences in data distributions, model architectures, and task objectives. Different tasks encourage models to learn distinct feature representations and thus possess different vulnerabilities. For instance, image–text retrieval focuses on global cross-modal similarity, whereas image captioning requires fine-grained visual understanding and linguistic generation. Consequently, UAPs that degrade retrieval features may not effectively disrupt captioning models.
Nevertheless, the proposed method achieves the strongest universal adversarial performance. This is attributed that it captures both local and global utilities of UAPs and increases data diversity to explore cross-modal interactions at multiple granularities. The future-aware momentum further alleviates overfitting, and the incorporation of text attacks adds complementary cross-modal disruption.
@@
| HRAimp w/o BM | 67.69 | 79.62 | 66.67 | 80.51 | 91.41 | 94.59 | 41.35 | 67.78 | 39.26 | 60.82 | 34.10 | 68.50 | 48.79 | 75.98 |
| HRAimp w/o ScMix | 72.54 | 84.70 | 72.03 | 85.73 | 95.34 | 96.91 | 50.06 | 78.51 | 47.36 | 72.42 | 33.30 | 66.10 | 49.53 | 76.67 |
| HRAimp w/o LUE | 72.80 | 83.22 | 69.73 | 84.97 | 94.72 | 96.13 | 52.88 | 79.16 | 43.68 | 73.42 | 33.99 | 68.15 | 47.42 | 74.76 |
| HRAimp w/o | 72.03 | 81.17 | 69.22 | 80.31 | 95.58 | 97.00 | 48.02 | 77.84 | 40.98 | 70.14 | 30.10 | 64.50 | 46.79 | 72.98 |
| HRAimp w/o | 72.29 | 83.22 | 69.99 | 84.91 | 95.09 | 95.36 | 48.90 | 76.32 | 38.42 | 67.65 | 29.83 | 61.55 | 45.37 | 69.29 |
| HRAimp w/o | 63.22 | 82.33 | 59.51 | 82.90 | 87.12 | 88.72 | 48.96 | 78.77 | 40.25 | 71.88 | 32.33 | 65.43 | 46.05 | 76.07 |
| HRAimp w/o | 62.84 | 74.55 | 58.49 | 75.71 | 94.23 | 96.84 | 42.09 | 68.56 | 40.47 | 63.92 | 30.06 | 63.83 | 42.68 | 73.48 |
@@ IV-C2 From visual grounding to image-text retrieval
We also evaluate transfer attacks from visual grounding to image–text retrieval, with results provided in Table V. The consistent improvements across models further demonstrate the effectiveness of the proposed method.
In summary, the strong performance on both cross-model and cross-task transfer scenarios verifies the effectiveness and generality of the proposed approach.
@@ IV-D1 Influence of Perturbation Budgets
The perturbation magnitude plays a crucial role in balancing attack effectiveness and imperceptibility. A small budget may fail to generate sufficiently disruptive perturbations, whereas a large budget can degrade visual quality and make the attack more easily detectable. Figure 2 and 3 summarize the results under varying perturbation magnitudes for image and text attacks respectively, which demonstrate our statement. Specifically, as the magnitude of UAPs increases, the performance of attacks increases. In addition, across all magnitudes, the proposed method consistently achieves the best performance, further verifying its superiority.
@@ IV-D2 Influence of Combination Weights in Future-Aware Momentum
We further analyze the role of past and future gradients by evaluating attack performance under different weighting schemes. We take HRAimp as an example, and the corresponding results are shown in Figure 4. From the figures, we can have the following observations. For past gradients, very small weights are insufficient to meaningfully regularize the current update, whereas excessively large weights overly constrain the optimization to previous directions and weaken the attack. For future gradients, small weights already provide beneficial guidance by anticipating upcoming trends, while large weights lead to overfitting.
@@ IV-D3 Influence of the Number of Future Steps
To investigate how future gradients contribute to improved transferability, we use HRAimp as an example and report the attack performance in Figure 5. The results show that a longer update horizon generally enhances performance. However, in some cases, an excessively long horizon begins to degrade it. For example, when attacking CLIP, extending the horizon from 1 to 2 steps improves transferability, whereas further increasing it from 2 to 4 steps leads to a slight performance drop. This degradation occurs because too many update steps cause the perturbation to overfit the source model’s gradients, thereby weakening its transferability to other models.
In terms of computational cost, on a workstation equipped with an NVIDIA L40s GPU, the training time of the proposed method with a number of steps of is , respectively. Overall, step = 2 provides the best trade-off between performance and efficiency, and the associated computational overhead remains acceptable.
@@ IV-E Ablation study
To verify the effectiveness of the proposed method, we conduct extensive ablation studies using HRAimp as an example. First, we evaluate the contribution of each component by removing them individually, resulting in HRAimp w/o LUE (local utility enhancement), HRAimp w/o ScMix, and HRAimp w/o FM (future-aware momentum). We also examine combinations of component removals, including HRAimp w/o , HRAimp w/o , HRAimp w/o , and HRAimp w/o . Second, we assess the roles of image and text attacks by disabling each separately, denoted as HRAimp w/o text attack and HRAimp w/o image attack. Third, we compare our approach with several existing augmentation or transformation strategies, including Random Crop (HRA w Crop), MixUp (HRA w MixUp), Block-wise Augmentation [36] (HRA w BWA), Translation [9] (HRA w TIA) and Rescale [24] (HRA w SIA). Fourth, we replace future-aware momentum with conventional momentum, denoted as (HRA w Mom).
Table VI summarizes all variant results under the image–text retrieval setting, using CLIP and Flickr30K as the source model and dataset, respectively. The results show that each proposed component contributes positively to attack performance and consistently outperforms corresponding alternatives. These findings further demonstrate the feasibility and importance of considering both the characteristics of UAPs and the optimization path.
@@ IV-F Visualization
In Figure 6, we showcase some examples of image-text retrieval under the proposed attack. From the figure, it is evident that the proposed method can effectively mislead the target model to return incorrect retrieved data. Additionally, we present some Grad-CAM [30] visualization examples in Figure 7, where it can be seen that the UAPs can significantly change the attention of the target model. These further confirm the effectiveness of the proposed method.
@@ V-A Conclusions
In this paper, we propose universal multimodal adversarial attacks that aim to efficiently explore the vulnerabilities of VLP models. To enhance adversarial transferability, we introduce a Hierarchical Refinement Attack that jointly explores the sample-agnostic, model-agnostic, and task-agnostic properties of UAPs. Our method performs refinement at both the sample level and the optimization level to produce more transferable perturbations. Extensive experiments across various VLP models, downstream tasks, and datasets demonstrate the effectiveness of the proposed attack.
@@ V-B Limitations and Future Work
Despite the significant improvement in universal adversarial transferability, the text attack implemented by inserting a uniform word across different sentences remains perceptible to humans. This is due to the discrete nature of text, even though we restrict the perturbation magnitude. In future work, we aim to develop more imperceptible text-based attacks. Additionally, adversarial transferability is still limited in certain settings, particularly under low perturbation budgets. We plan to further investigate the underlying mechanisms of model perception and explore more effective and transferable universal attack strategies.

### v2

@@ IV-B2 Effects of multimodal attacks
Incorporating text attacks further enhances overall performance, as reflected by the gains of C-PGC and the proposed HRA over other methods. Replacing the most salient words provides additional improvement, evident from the superior results of HRA compared with HRA. Interestingly, although C-PGC also replaces important words, HRA still outperforms it in many cases, highlighting the effectiveness of directly discovering substitution words rather than learning universal embeddings and mapping them back to tokens. Moreover, our method does not rely on a predefined word library, demonstrating its practicality. While identifying the most important words incurs additional computational overhead (HRAimp), HRAran avoids this cost.
@@ IV-B3 Effects of learning objectives
Learning objectives also affect transferability. Models trained with similar objectives generally exhibit stronger adversarial transfer, as observed across different CLIP variants and other architectures. For example, UAPs generated on CLIP transfer more effectively to CLIP than to ALBEF or TCL.
@@ IV-C1 From image-text retrieval to other tasks
We first examine transfer attacks from image–text retrieval to visual grounding and image captioning, with results presented in Tables II and III, respectively. For both experiments, we generate UAPs using CLIP and ALBEF as the source models and Flickr30K as the source dataset. For visual grounding, we use ALBEF and RefCOCO+ as the target model and dataset. For image captioning, we use BLIP and MSCOCO as the target model and dataset.
Across both tasks, all methods exhibit notably reduced transferability. This is expected due to substantial differences in data distributions, model architectures, and task objectives. Different tasks encourage models to learn distinct feature representations and thus possess different vulnerabilities. For instance, image–text retrieval focuses on global cross-modal similarity, whereas image captioning requires fine-grained visual understanding and linguistic generation. Consequently, UAPs that degrade retrieval features may not effectively disrupt captioning models. Nevertheless, the proposed method achieves the strongest universal adversarial performance. This is attributed to its capability to prevent overfitting by considering past and future update directions and the incorporation of text attacks via hierarchical modeling of textual importance.
@@
| HRAimp w/o FM | 67.69 | 79.62 | 66.67 | 80.51 | 91.41 | 94.59 | 41.35 | 67.78 | 39.26 | 60.82 | 34.10 | 68.50 | 48.79 | 75.98 |
@@ IV-C2 From visual grounding to image-text retrieval
We also evaluate transfer attacks from visual grounding to image–text retrieval, with results reported in Table IV. The consistent improvements across models further confirm the effectiveness of the proposed method.
In summary, the strong performance across cross-model and cross-task transfers on different datasets validates the effectiveness and transferability of our approach.
@@ IV-D1 Influence of Perturbation Budgets
The perturbation magnitude plays a critical role in balancing attack effectiveness and imperceptibility. Figures 2 and 3 present the results of image and text attacks under varying magnitudes. From the figures, it can be observed that attack performance generally improves as the magnitude of UAPs increases. However, excessively large perturbations can degrade visual quality and make attacks more detectable. Importantly, across all magnitudes, the proposed method consistently outperforms competing approaches, further highlighting its superiority.
@@ IV-D2 Influence of Combination Weights in Future-Aware Momentum
We further analyze the contributions of past and future gradients by evaluating attack performance under different weighting schemes, using HRAimp as an example. The corresponding results are shown in Figure 4, from which critical observations can be made. For past gradients, very small weights fail to sufficiently regularize the current update, whereas excessively large weights overly constrain the optimization to historical directions, reducing attack effectiveness. For future gradients, even small weights provide useful guidance by anticipating upcoming trends, while large weights can also lead to overfitting to the future optimization trend.
@@ IV-D3 Influence of the Number of Future Steps
To investigate how future gradients contribute to improved transferability, we use HRAimp as an example and report the attack performance in Figure 5. The results show that a longer update horizon generally enhances performance. However, in some cases, an excessively long horizon begins to degrade it. For example, when attacking CLIPViT-B/16, extending the horizon from 1 to 2 steps improves transferability, whereas further increasing it from 2 to 4 steps leads to a slight performance drop. This degradation occurs because too many update steps cause the perturbation to overfit the source model’s gradients, thereby weakening its transferability to other models.
In terms of computational cost for incorporating future update information, on a workstation equipped with an NVIDIA L40s GPU, the training time of the proposed method for , , , , and steps is , , , , and seconds, respectively. Overall, using steps provides the best trade-off between performance and efficiency, with an acceptable computational overhead.
@@ IV-E Ablation study
To verify the effectiveness of each component of the proposed method, we conduct extensive ablation studies using HRA as an example. First, we evaluate the contribution of individual components by removing them separately, resulting in HRA w/o FM (future-aware momentum), HRA w/o Text attack, and HRA w/o Image attack. Second, we assess our approach with several existing augmentation strategies, including Random Crop (HRA w Crop), MixUp (HRA w MixUp), Block-wise Augmentation [33] (HRA w BWA), Translation [6] (HRA w TIA), and Rescale [21] (HRA w SIA). Third, we replace future-aware momentum with conventional momentum, denoted as HRA w Mom.
Table V summarizes the results of all variants under the image–text retrieval setting, using CLIP and Flickr30K as the source model and dataset, respectively. The results indicate that each proposed component contributes positively to attack performance and consistently outperforms the corresponding alternatives. In addition, the proposed method can be combined with different augmentation strategies to further enhance performance. These findings further validate the effectiveness of the proposed method.
@@ IV-F Visualization
Figure 6 presents examples of image–text retrieval under the proposed attack. The results show that the method can effectively mislead the target model into returning incorrect retrievals while remaining visually imperceptible. dditionally, we present some Grad-CAM [27] visualization examples in Figure 7, where it can be seen that the UAPs can significantly change the attention of the target model. These further confirm the effectiveness of the proposed approach.
@@ V-A Conclusions
In this paper, we propose a universal multimodal adversarial attack framework that systematically explores the vulnerabilities of VLP models and targets the specific characteristics of attacks in each modality. To improve adversarial transferability, we introduce a Hierarchical Refinement Attack (HRA) that hierarchically regularizes the optimization trajectory for image UAPs, mitigating overfitting caused by convergence to local optima. For the text modality, we design a practical and effective UAP learning strategy that accounts for both intra- and inter-sentence importance, ensuring globally influential word replacements. Extensive experiments across diverse VLP models, downstream tasks, and datasets demonstrate the effectiveness and transferability of the proposed attacks.
@@ V-B Limitations and Future Work
Despite the significant improvements in universal adversarial transferability, the text attack that inserts a uniform word across different sentences remains perceptible to human readers. This limitation stems from the discrete nature of text, even when the perturbation magnitude is strictly constrained. In future work, we aim to develop more imperceptible text-based attack strategies. Moreover, adversarial transferability remains limited in certain scenarios, particularly under low perturbation budgets. We plan to further investigate the underlying mechanisms of model perception and explore more effective and transferable universal attack methods.